# Renormalizable Spectral-Shell Dynamics as the Origin of Neural Scaling Laws

### v3

@@ Abstract
To extract macroscopic behavior, we introduce a logarithmic spectral-shell coarse-graining and track quadratic error energy across shells. Microscopic interactions within each shell cancel identically at the energy level, so shell energies evolve only through dissipation and external inter-shell interactions. We formalize this via a renormalizable shell-dynamics assumption, under which cumulative microscopic effects reduce to a controlled net flux across shell boundaries.
Assuming an effective power-law spectral transport in a relevant resolution range, the shell dynamics admits a self-similar solution with a moving resolution frontier and explicit scaling exponents. This framework explains neural scaling laws and double descent, and unifies lazy (NTK-like) training and feature learning as two limits of the same spectral-shell dynamics.
@@ 4.4 PDE-approximability across shells (energy continuum)
$E_{\alpha}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}\varepsilon(\lambda,t)\,d\lambda,\qquad\mathcal{L}(t)=\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda.$
@@
This is a local continuum approximation over that range; we do not require a global continuous-spectrum limit. As in nonequilibrium statistical physics forster1977large; spohn2012large, we do not assume the existence of an exact continuum limit; rather, renormalizability refers to the existence of a closed effective description after coarse-graining. We stress again that our continuum equations do not assume the existence of a genuine continuous-spectrum limit. They represent a local approximation of the discrete shell-flux difference when shells are sufficiently dense over a dynamically relevant resolution range. Formally, this amounts to approximating the discrete net-flux difference $J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)$ by a first-order finite-difference representation of $\partial_{\lambda}J$ over a dense but finite shell range.
4.5 Assumption: effective power-law transport (energy-flux interface)
Assumption 1 (Effective power-law transport (energy form)).
There exists a dynamically relevant resolution range $\lambda\in[\lambda_{\min},\lambda_{\max}]$ and a time-dependent amplitude $c(t)$ such that the renormalized energy flux can be expressed as
$J(\lambda,t)=v(\lambda,t)\,\varepsilon(\lambda,t),\qquad v(\lambda,t)\sim-c(t)\lambda^{b},\quad b>0,$
up to subleading corrections consistent with Definition 4.1. Equivalently, in the same range the coarse-grained energy dynamics admit an effective transport–dissipation form
$\partial_{t}\varepsilon(\lambda,t)+\partial_{\lambda}\!\big(v(\lambda,t)\varepsilon(\lambda,t)\big)=-2\lambda\,\varepsilon(\lambda,t),$
to leading order (the factor $2$ reflects quadratic energy dissipation).
Remark 4 (Scope).
Assumption 1 summarizes the outcome of coarse-graining rather than a microscopic property of $M(t)$. Providing operator-level conditions ensuring renormalizability and effective power-law transport is deferred to companion work.
5 Renormalized Energy-Shell Dynamics and the GRSD Regime
In Sections 34, we derived the exact modewise dynamics, introduced logarithmic spectral shells, and identified the quadratic shell energy $E_{\alpha}(t)=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2}$ as the canonical conserved quantity for shell-internal coupling. We now formulate the macroscopic regime of interest: a high-resolution tail where the inter-shell exchange is renormalizable and can be summarized by a controlled net-flux across shell boundaries.
5.1 The closed shell-level bookkeeping
Recall the exact shell balance law (Section 3.4):
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t),\qquad D_{\alpha}(t):=\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2},$ (4)
where $\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t)$ is the inter-shell quadratic-energy exchange. In particular,
$\sum_{\alpha}\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t)=0,\qquad\frac{d}{dt}\Big(\sum_{\alpha}E_{\alpha}(t)\Big)=-\sum_{\alpha}D_{\alpha}(t)\leq 0.$
Thus, only dissipation decreases the total loss, while coupling redistributes energy across shells.
5.2 Renormalizable shell dynamics (energy form)
We work in the macroscopic closure regime postulated in Definition 4.1. Concretely, for sufficiently large $\alpha$ we assume the inter-shell exchange can be summarized as a controlled net-flux across the shell boundary $\alpha$ plus a negligible remainder:
$\frac{d}{dt}E_{\alpha}(t)=-D_{\alpha}(t)+\big(J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)\big)+\mathcal{R}_{\alpha}(t),\qquad|\mathcal{R}_{\alpha}(t)|\ll D_{\alpha}(t).$ (5)
This is a shell-level statement: it does not require a continuum limit, and it does not assume microscopic locality of coupling. It only asserts that after coarse-graining, the net exchange across shell boundaries is well-defined and controlled.
5.3 Continuum approximation (energy density)
Over a resolution range where shells are sufficiently fine, we may introduce an energy density $\varepsilon(\lambda,t)$ such that
$E_{\alpha}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}\varepsilon(\lambda,t)\,d\lambda,\qquad\mathcal{L}(t)=\sum_{\alpha}E_{\alpha}(t)\approx\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda.$
In this regime, the discrete net-flux difference $J_{\leq\alpha}-J_{\leq\alpha-1}$ may be approximated by a transport divergence $\partial_{\lambda}J(\lambda,t)$. Here “sufficiently fine” refers only to the scale separation between neighboring shells relative to the variation of the cumulative flux, not to any limiting process in which the shell spacing tends to zero.
5.4 Effective power-law transport (energy-flux interface)
We now specialize to the single effective interface used later (Assumption 1):
$J(\lambda,t)=v(\lambda,t)\,\varepsilon(\lambda,t),\qquad v(\lambda,t)\sim-c(t)\lambda^{b},\quad b>0,$ (6)
which yields, to leading order, the energy transport–dissipation equation
@@ 5.5 RMS amplitude notation
$\mathcal{L}(t)=\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda=\frac{1}{2}\int_{0}^{\infty}g(\lambda,t)^{2}\,d\lambda,$
@@
5.6 Summary
The key structural takeaway is:
All scaling-law consequences in later sections are derived from the renormalized boundary-flux interface (5)–(6) in a high-resolution range.
6 Self-Similar GRSD Tail Under Power-Law Drift
In this section we analyze the effective GRSD regime implied by the energy transport interface (7) with power-law drift $v(\lambda,t)\sim-c(t)\lambda^{b}$. We emphasize that the discrete shell bookkeeping is primary; the continuum calculation below is an approximation valid over a dense shell range. Equivalently, all results in this section can be interpreted as asymptotic solutions of the discrete shell-energy dynamics, with the PDE serving only as a convenient analytic representation.
6.1 Energy transport equation and characteristics
We start from the leading-order energy PDE
$\partial_{t}\varepsilon+\partial_{\lambda}(v\varepsilon)=-2\lambda\,\varepsilon,\qquad v(\lambda,t)\sim-c(t)\lambda^{b}.$
Expanding the transport term gives
$\partial_{t}\varepsilon+v\,\partial_{\lambda}\varepsilon+(\partial_{\lambda}v)\,\varepsilon=-2\lambda\,\varepsilon.$ (9)
Along characteristics $\lambda(t)$ satisfying
$\frac{d\lambda}{dt}=v(\lambda,t)\sim-c(t)\lambda^{b},$ (10)
the energy density evolves by
$\frac{d}{dt}\log\varepsilon(\lambda(t),t)=-\partial_{\lambda}v(\lambda(t),t)-2\lambda(t).$ (11)
Define the effective training time
$\tau(t):=\int_{0}^{t}c(s)\,ds.$ (12)
For $b\neq 1$, the characteristic flow integrates to
$\lambda(t)^{1-b}=\lambda_{0}^{1-b}+(b-1)\tau(t),\qquad(b\neq 1).$ (13)
In the supercritical regime $b>1$, characteristics move from high to lower resolution progressively, enabling a resolution cascade over time.
6.2 Universal tail factor from transport Jacobian
Using $\partial_{\lambda}v(\lambda,t)\sim-bc(t)\lambda^{b-1}$, equation (11) becomes
$\frac{d}{dt}\log\varepsilon(\lambda(t),t)=bc(t)\lambda(t)^{b-1}-2\lambda(t).$
The first term is purely geometric: it comes from the Jacobian of the spectral flow (compression/expansion in $\lambda$), while the second term is physical dissipation.
Ignoring dissipation temporarily, the transport-only equation $\partial_{t}\varepsilon+\partial_{\lambda}(v\varepsilon)=0$ yields the universal scaling
$\varepsilon(\lambda,t)\propto\lambda^{-b}.$ (14)
Equivalently, the RMS amplitude density satisfies the transport-only tail
$g(\lambda,t)\propto\lambda^{-b/2}.$ (15)
This tail exponent is fixed by the power-law drift and does not depend on the microscopic details of $M(t)$, only on the renormalized transport interface.
6.3 Dissipation and the resolution frontier
We now account for dissipation. Along a characteristic, define the dissipation action
$\Phi(\lambda,t):=\int_{0}^{t}2\lambda(s)\,ds,$ (16)
where $\lambda(s)=\lambda(s;\lambda_{0})$ follows (10). In the supercritical regime $b>1$, inserting the asymptotic behavior $\lambda(t)\sim\big((b-1)\tau(t)\big)^{-1/(b-1)}$ into (16) yields the scaling
$\Phi(\lambda,t)\;\asymp\;K\,\lambda^{\,b-1}\,\tau(t),$ (17)
for a constant $K>0$ (the exact prefactor is not important for scaling).
This produces a sharp cutoff in the energy density: modes with $\lambda^{b-1}\tau(t)\gg 1$ are exponentially suppressed. Hence the dynamics localize around a resolution frontier
$\lambda^{\ast}(t)\;\asymp\;\tau(t)^{-1/(b-1)}.$ (18)
6.4 GRSD tail form (energy and amplitude)
Combining the universal transport tail (14) with the dissipation action (17), we obtain the GRSD tail template in energy form:
$\boxed{\varepsilon(\lambda,t)=A_{\varepsilon}(t)\,\lambda^{-b}\,\exp\!\big(-K\,\lambda^{\,b-1}\tau(t)\big),}$ (19)
where $A_{\varepsilon}(t)$ is a slowly varying (typically polynomial) amplitude determined by matching to lower-resolution shells / initial data.
Equivalently, in RMS amplitude notation $g=\sqrt{2\varepsilon}$,
$\boxed{g(\lambda,t)=A_{g}(t)\,\lambda^{-b/2}\,\exp\!\Big(-\tfrac{1}{2}K\,\lambda^{\,b-1}\tau(t)\Big),\qquad A_{g}(t):=\sqrt{2A_{\varepsilon}(t)}.}$ (20)
6.5 Remarks on regimes $b\leq 1$
The critical case $b=1$ corresponds to exponential-in-$\tau$ characteristics and does not produce a progressive resolution frontier. The subcritical case $0<b<1$ reaches the resolution floor in finite $\tau$ and therefore cannot sustain asymptotic data scaling. Since modern scalable training empirically exhibits progressive refinement, we focus on $b>1$ in the main text and discuss $b\leq 1$ in Appendix B.
6.6 Summary
Under the renormalized shell-flux interface and effective power-law drift, the high-resolution tail admits a universal GRSD form: a power-law transport tail multiplied by an exponential dissipation cutoff, with a resolution frontier $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$.
7 Scaling Laws and the Geometry of Double Descent
We now derive scaling-law behavior from the GRSD template. The correct global loss is the total quadratic error energy
$\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\sum_{\alpha}E_{\alpha}(t)\approx\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda=\frac{1}{2}\int_{0}^{\infty}g(\lambda,t)^{2}\,d\lambda.$
Thus the scaling analysis should be performed on $\varepsilon(\lambda,t)$ (or equivalently $g^{2}$).
7.1 Loss scaling from the GRSD tail
Insert the energy-form GRSD template (19):
$\varepsilon(\lambda,t)=A_{\varepsilon}(t)\lambda^{-b}\exp\!\big(-K\lambda^{b-1}\tau(t)\big).$
For large $\tau(t)$, the integral is dominated near the frontier $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$. A standard Laplace-type estimate yields
$\mathcal{L}(t)\;\asymp\;A_{\varepsilon}(t)\,\big(\lambda^{\ast}(t)\big)^{1-b}.$ (21)
Using $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$ gives
$\mathcal{L}(t)\;\asymp\;A_{\varepsilon}(t)\,\tau(t)^{\frac{b-1}{b-1}\cdot\frac{b-1-(1)}{\,b-1\,}}=A_{\varepsilon}(t)\,\tau(t)^{\frac{b-1-(1)}{\,b-1\,}}=A_{\varepsilon}(t)\,\tau(t)^{-\frac{b-2}{\,b-1\,}}.$ (22)
Hence, once $A_{\varepsilon}(t)$ is controlled (typically polynomial/slowly varying), the dominant exponent is determined by $b$.
Remark 5 (Where $A_{\varepsilon}(t)$ comes from).
$A_{\varepsilon}(t)$ is not an arbitrary free function: it is fixed by matching the high-resolution tail to the lower-resolution part of the spectrum and by conservation/dissipation constraints at the shell level. In particular, $A_{\varepsilon}(t)$ encodes how much energy is injected into the tail by inter-shell transport at a given effective time. Our framework isolates the universal tail shape; the prefactor can be estimated empirically from spectral measurements.
7.2 Double Descent from In-Domain Finite-Sample Spectral Mismatch
Training dynamics and monotone training loss.
Throughout training, the error evolves according to
Measured in the training metric, the quadratic loss
is strictly non-increasing:
$\frac{d}{dt}\mathcal{L}_{\rm tr}(t)=-\langle e_{t},M(t)e_{t}\rangle\leq 0.$
In-domain train/test mismatch as a finite-sample effect.
$\Delta:=P_{\rm te}-P_{\rm tr}.$
The test loss can be written as
Shell amplification via geometric growth of $A(t)$.
The effect of this mismatch becomes dynamically visible through the shell structure of the error. Under the renormalizable shell-dynamics assumption, the high-resolution tail of the error admits the generalized resolution spectral form
$g(\lambda,t)\;\approx\;A(t)\,\lambda^{-b}\,e^{-\phi(\lambda,t)},\qquad b>1,$
over a dynamically relevant range of resolutions. As shown in Section LABEL:sec:weak_coupling, the geometric prefactor $A(t)$ grows polynomially during an initial training phase, reflecting compression of spectral coordinates induced by transport across resolution shells.
Since the mismatch operator $\Delta$ is concentrated in high-resolution shells, the correction term satisfies a lower bound of the form
$\langle e_{t},\Delta e_{t}\rangle\;\gtrsim\;\sum_{\alpha\geq\alpha_{0}}\delta_{\alpha}\,E_{\alpha}(t),\qquad\delta_{\alpha}>0,$
for some tail index $\alpha_{0}$. In the early-time regime where dissipation is not yet active on the tail, the shell energies scale as
$\sum_{\alpha\geq\alpha_{0}}E_{\alpha}(t)\;\asymp\;A(t)^{2}\int_{\lambda_{\alpha_{0}}}^{\infty}\lambda^{-2b}\,d\lambda,$
so that the test correction grows proportionally to $A(t)^{2}$. As a result, the growth of $A(t)$ can temporarily dominate the monotone decay of $\mathcal{L}_{\rm tr}(t)$, forcing the test loss $\mathcal{L}_{\rm te}(t)$ to increase even though training loss continues to decrease.
Dissipation, resolution frontier, and eventual decay.
At later times, dissipation encoded by the factor $e^{-\phi(\lambda,t)}$ suppresses the spectral tail and localizes the error near the resolution frontier $\lambda^{*}(t)$. Both training and test losses then decay to zero, and the finite-sample mismatch term becomes negligible. The test loss therefore exhibits a rise–then–fall profile: an initial increase driven by shell-level geometric amplification, followed by universal decay once dissipation dominates.
Interpretation.
Double descent thus emerges as a generic consequence of three ingredients: (i) exact monotonic dissipation in the training metric, (ii) unavoidable finite-sample spectral mismatch in high-resolution modes, and (iii) geometric amplification of tail energy induced by spectral-shell transport. This mechanism operates entirely in-domain and does not rely on distribution shift, overparameterization heuristics, or ad hoc regularization. Instead, it reflects a structural interaction between sampling noise and the renormalizable shell dynamics governing neural training.
@@ 7.4 Summary
In the GRSD regime ($b>1$), the renormalized energy-shell dynamics yield: (i) a universal high-resolution tail shape (19), (ii) a resolution frontier $\lambda^{\ast}(t)\asymp\tau(t)^{-1/(b-1)}$, and (iii) scaling-law decay of the total loss controlled by the drift exponent $b$ and the effective time $\tau(t)$. Under standard optimization assumptions, $\tau(t)$ grows at most polynomially (Appendix C).

### v4

@@ Abstract
To extract macroscopic behavior, we introduce a logarithmic spectral-shell coarse-graining and track quadratic error energy across shells. Microscopic interactions within each shell cancel identically at the energy level, so shell energies evolve only through dissipation and external inter-shell interactions. We formalize this via a renormalizable shell-dynamics assumption, under which cumulative microscopic effects reduce to a controlled net flux across shell boundaries. This shell-form equation unifies lazy-regime training and feature learning as two limits of the same spectral-shell dynamics.
Assuming an effective power-law spectral transport in a relevant resolution range, the shell dynamics admits a self-similar solution with a moving resolution frontier and explicit scaling exponents, which explains neural scaling laws and double descent.
@@ 4.4 PDE-approximability across shells (energy continuum)
$E_{\alpha}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}\varepsilon(\lambda,t)\,d\lambda,\qquad J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)\approx\int_{\lambda_{\alpha}}^{\lambda_{\alpha+1}}J(\lambda,t)\,d\lambda,$
@@
This is a local continuum approximation over that range; we do not require a global continuous-spectrum limit. As in nonequilibrium statistical physics forster1977large; spohn2012large, we do not assume the existence of an exact continuum limit; rather, renormalizability refers to the existence of a closed effective description after coarse-graining. We stress again that our continuum equations do not assume the existence of a genuine continuous-spectrum limit. Under this approximation we acquire the expression of the total loss function:
$\mathcal{L}(t)\approx\int_{\lambda_{\min}}^{\lambda_{\max}}\varepsilon(\lambda,t)\,d\lambda.$
They represent a local approximation of the discrete shell-flux difference when shells are sufficiently dense over a dynamically relevant resolution range. Formally, this amounts to approximating the discrete net-flux difference $J_{\leq\alpha}(t)-J_{\leq\alpha-1}(t)$ by a first-order finite-difference representation of $\partial_{\lambda}J$ over a dense but finite shell range. This energy continumm yields, to leading order, the energy transport–dissipation equation
$\partial_{t}\varepsilon(\lambda,t)+\partial_{\lambda}J(\lambda,t)=-2\lambda\,\varepsilon(\lambda,t),$ (4)
@@ 4.5 RMS amplitude notation
$\mathcal{L}(t)\approx\int_{\lambda_{\min}}^{\lambda_{\max}}\varepsilon(\lambda,t)\,d\lambda=\int_{\lambda_{\min}}^{\lambda_{\max}}g(\lambda,t)^{2}\,d\lambda,$
@@ 4.7 Summary
The key structural takeaway is:
All scaling-law consequences in later sections are derived from the renormalized boundary-flux interface (4.1)–(4) in a high-resolution range.
5 Scale-Free GRSD Under Power-Law–Compatible Transport
In this section we analyze the effective GRSD regime arising in an intermediate scale-free training window. Empirically, modern deep models exhibit extended intervals $t_{0}<t<T$ during which macroscopic spectral observables evolve as straight lines in log–log coordinates, indicating the absence of intrinsic scales.
Within such a window, the effective spectral transport must itself be approximately scale-free. While the most general scale-free drift takes the form $v(\lambda,t)=\phi(\lambda/t)$, requiring pure power-law temporal scaling of observables rules out generic $\phi$ and selects the linear scale-free transport
$v(\lambda,t)=c_{0}\,\frac{\lambda}{t},\qquad t_{0}<t<T,$ (6)
up to subleading corrections. In the remainder of this section we take (6) as an effective model assumption and solve the resulting GRSD dynamics explicitly.
Why scale-free observables rigidly constrain the drift.
We emphasize that the power-law form
$v(\lambda,t)=c_{0}\,\frac{\lambda}{t}$
is not introduced as an arbitrary modeling choice. Rather, it is rigidly selected by the requirement that macroscopic spectral observables remain scale-free over an extended training window.
Empirically, in this regime, shell-resolved quantities such as the spectral energy density, cumulative shell energies, and total loss evolve as straight lines in log–log coordinates over multiple decades of time and resolution. This behavior implies not merely scale invariance at a fixed time, but a joint scale covariance between the spectral coordinate $\lambda$ and the training time $t$.
At the level of the coarse-grained shell dynamics, this covariance severely restricts the admissible form of the effective transport velocity. While the most general scale-free drift compatible with dimensional analysis can be written as $v(\lambda,t)=\phi(\lambda/t)$, generic choices of $\phi$ induce nonlinear reparameterizations of the spectral coordinate and therefore produce curvature in log–log observables. Requiring that these observables remain asymptotically affine in log–log space selects the linear form $\phi(u)=c_{0}u$, yielding $v(\lambda,t)=c_{0}\lambda/t$ up to subleading corrections.
Lagrangian spectral coordinates as the renormalized frame.
The emergence of the Lagrangian spectral coordinate $u$ should not be viewed as a technical device introduced solely to solve the transport equation. Instead, it represents the natural renormalized variable induced by the shell-level flux interface.
At the discrete shell level, renormalizability asserts that the cumulative effect of inter-shell interactions can be summarized by a controlled boundary flux. In the continuum approximation, this structure manifests as a transport term that can be absorbed into a time-dependent reparameterization of the spectral coordinate. The characteristic flow (5) implements precisely this absorption.
In the co-moving coordinate $u$, the dynamics reduce to pure dissipation. All nontrivial macroscopic structure—including scaling laws and apparent emergence— is therefore encoded in how this co-moving frame maps back to stationary spectral coordinates. The simplicity of the resulting evolution is not accidental, but a direct consequence of the renormalizable shell-dynamics assumption.
5.1 Characteristic flow and scale-free reparameterization
Starting from the leading-order GRSD equation
with drift (6), the characteristic curves satisfy
$\frac{d\lambda}{dt}=c_{0}\,\frac{\lambda}{t}.$ (8)
Integrating yields the exact scale-free flow
$\lambda(t;u)=u\Big(\frac{t}{t_{0}}\Big)^{c_{0}},\qquad u=\lambda\Big(\frac{t}{t_{0}}\Big)^{-c_{0}},$ (9)
where $u$ defines a stationary (Lagrangian) spectral coordinate. This flow is strictly power-law and affine in log–log coordinates.
5.2 Energy density along characteristics
Along a characteristic $\lambda(t;u)$, the energy density evolves as
$\frac{d}{dt}\log\varepsilon(\lambda(t),t)=-2\lambda(t).$ (10)
Substituting (9) and integrating gives
$\varepsilon(u,t)=\varepsilon_{0}(u)\,\exp\!\Big[-\frac{2u}{1+c_{0}}\Big(\Big(\frac{t}{t_{0}}\Big)^{1+c_{0}}-1\Big)\Big],\qquad c_{0}\neq-1.$ (11)
which is aligned with the spectral evolution conjecture proposed in zhang2025generalized:
$f^{*}-f_{\lambda}(t)=w_{\lambda}\exp(g(\lambda,t))$ (12)
where $g(\lambda,t)\propto\lambda^{a(\beta)}t^{b(\beta)}$ and thus $w_{\lambda}^{2}$ correspond to the starting error on mode with eigenvalue of $\lambda$. This conjecture is helpful for understanding the scaling law for both loss evolution and model compression. It further predicted the existence of learning frontier and defines model density, which is observed to evolve according to Xiao2024DensingLaw.
Rewriting (11) in Eulerian coordinates using $u=\lambda(t/t_{0})^{-c_{0}}$ yields
$\boxed{\varepsilon(\lambda,t)=\varepsilon_{0}\!\big(\lambda\,t^{-c_{0}}\big)\,\exp\!\Big(-\frac{2\lambda}{1+c_{0}}\,t\Big)}.$ (13)
The exponential factor represents physical dissipation, while the prefactor encodes pure scale-free transport. The competition between transport and dissipation is governed entirely by the combination $1+c_{0}$. This expression constitutes the leading-order GRSD solution in the scale-free window.
The total quadratic loss is
$\mathcal{L}(t)=\int_{0}^{\infty}\varepsilon(\lambda,t)\,d\lambda.$
Substituting (13) and performing a standard moving-front estimate yields power-law decay
$\mathcal{L}(t)\sim t^{-(1+c_{0})(1-\alpha)},$
for initial spectra $\varepsilon_{0}(u)\sim u^{-\alpha}$.
Transport–dissipation competition and spectral front formation.
The explicit solution reveals a universal structural feature: the dynamics are governed by a competition between scale-free transport and scale-dependent dissipation. Transport alone redistributes error mass across resolutions without loss, while dissipation selectively suppresses high-$\lambda$ components.
Their interplay generates a moving spectral front separating a learned region from an unresolved tail. Importantly, this front is not an artifact of a particular initialization, architecture, or optimizer choice. It is a robust consequence of renormalizable spectral transport under power-law–compatible drift.
In the next section, we show that when training and evaluation are performed on finite samples drawn from the same underlying distribution, this moving front interacts with unavoidable high-resolution spectral mismatch. The resulting transient amplification in certain spectral bands provides a natural and entirely in-domain explanation of the double-descent phenomenon.
6 Double Descent from In-Domain Finite-Sample Spectral Mismatch
6.1 Training dynamics and monotone training loss.
Throughout training, the error evolves according to
Measured in the training metric, the quadratic loss
is strictly non-increasing:
$\frac{d}{dt}\mathcal{L}_{\rm tr}(t)=-\langle e_{t},M(t)e_{t}\rangle\leq 0.$
6.2 In-domain train/test mismatch as a finite-sample effect.
$\Delta:=P_{\rm te}-P_{\rm tr}.$
The test loss can be written as
Double descent as a generic finite-sample effect.
We stress that the non-monotonic behavior described here does not rely on distribution shift, adversarial sampling, or overparameterization per se. It arises generically whenever the population operator exhibits a slowly decaying spectral tail and learning proceeds through a transport-dominated regime.
From this perspective, double descent is not a pathological deviation from monotone learning, but a predictable finite-sample manifestation of renormalizable spectral transport. As model capacity or dataset size increases, the location of the mismatch-dominated spectral region shifts, modifying the timing and magnitude of the transient amplification without eliminating the underlying mechanism.
6.3 Transport-induced amplitude reweighting
Starting from the GRSD solution (13), the spectral error density admits the exact factorization
$\varepsilon(\lambda,t)=A(\lambda,t)\,\exp\!\Big(-\frac{2\lambda}{1+c_{0}}\,t\Big),\qquad A(\lambda,t):=\varepsilon_{0}(\lambda t^{-c_{0}}).$ (14)
The exponential factor represents pure dissipation and is strictly decreasing in time. All non-monotonic behavior at fixed $\lambda$ is therefore encoded in the amplitude function $A(\lambda,t)$, which arises solely from scale-free transport.
Intuitive picture.
Intuitively, scale-free transport continuously transfers error mass from higher-$\lambda$ modes, which are learned and dissipated rapidly, toward lower-$\lambda$ modes. As a result, even as high-resolution errors decay, part of this error is temporarily deposited onto lower-resolution components, raising their spectral error density.
At the same time, dissipation at scale $\lambda$ is proportional to $\lambda\,\varepsilon(\lambda,t)$, and is therefore weaker at small $\lambda$. The mismatch between fast transport from high $\lambda$ and delayed dissipation at low $\lambda$ naturally produces a transient amplification before eventual decay.
Differentiating the amplitude function at fixed $\lambda$ yields
$\partial_{t}A(\lambda,t)=-\,c_{0}\,\lambda\,t^{-c_{0}-1}\,\varepsilon_{0}^{\prime}(\lambda t^{-c_{0}}).$ (15)
Hence, whenever the initial spectrum is decreasing, $\varepsilon_{0}^{\prime}(u)<0$, and $c_{0}>0$, the amplitude grows in time:
$\partial_{t}A(\lambda,t)>0.$
Consequently, the spectral error density $\varepsilon(\lambda,t)$ can increase transiently at fixed $\lambda$, despite overall dissipation and monotone decay of the training loss.
6.4 Double descent from transport-induced amplification in the mismatch window
We now formalize how transport-induced amplitude growth leads to double descent when combined with finite-sample spectral mismatch. The key point is that, although the training loss is strictly monotone, the test loss probes a different quadratic form that selectively weights high-resolution spectral components.
Spectral localization of finite-sample mismatch.
Recall that the test loss can be written as
$L_{\mathrm{te}}(t)=L_{\mathrm{tr}}(t)+\frac{1}{2}\langle e_{t},\Delta e_{t}\rangle,\qquad\Delta:=P_{\mathrm{te}}-P_{\mathrm{tr}},$
where $\Delta$ is the mismatch operator. Classical results on empirical covariance and integral operators imply that, for continuous data distributions and infinite-rank population operators, $\Delta$ is generically supported on high-resolution modes, corresponding to small eigenvalues $\lambda$ of the population operator.
At the coarse-grained level, this implies that there exists a resolution interval
$\lambda\in[\lambda_{-},\lambda_{+}],$
which we refer to as the mismatch window, such that the dominant contribution to $\langle e_{t},\Delta e_{t}\rangle$ comes from spectral components in this range. Outside this window, either the modes are well-aligned between train and test or their contribution is suppressed by dissipation.
Mismatch-weighted test loss and non-monotonicity.
The mismatch contribution to the test loss can be approximated, to leading order, by restricting to the mismatch window:
$\langle e_{t},\Delta e_{t}\rangle\;\approx\;\int_{\lambda_{-}}^{\lambda_{+}}w(\lambda)\,\varepsilon(\lambda,t)\,d\lambda,$
where $w(\lambda)\geq 0$ denotes the coarse-grained spectral weight induced by the finite-sample mismatch operator $\Delta=P_{\mathrm{te}}-P_{\mathrm{tr}}$ in the instantaneous spectral basis of $M(t)$, so that $w(\lambda)\,d\lambda$ represents the leading-order contribution of modes with eigenvalues in $\lambda,\lambda+d\lambda]$ to the quadratic form $\langle e_{t},\Delta e_{t}\rangle$. Formally, $w(\lambda)$ is defined by coarse-graining the diagonal spectral density of the mismatch operator $\Delta$ in the instantaneous eigenbasis of $M(t)$, i.e.
$w(\lambda)\;:=\;\sum_{u:\,\lambda_{u}(t)\in[\lambda,\lambda+d\lambda)}\langle\phi_{u}(t),\Delta\,\phi_{u}(t)\rangle\,/\,d\lambda,$
so that $\langle e_{t},\Delta e_{t}\rangle\approx\int w(\lambda)\,\varepsilon(\lambda,t)\,d\lambda$ to leading order.
Substituting the factorized form of $\varepsilon$ gives
$\int_{\lambda_{-}}^{\lambda_{+}}w(\lambda)\,A(\lambda,t)\,\exp\!\left(-\frac{2\lambda}{1+c_{0}}\,t\right)\,d\lambda.$
If the mismatch window $\lambda_{-},\lambda_{+}]$ intersects the region of spectral coordinates for which $\partial_{t}A(\lambda,t)>0$, then the integrand exhibits a competition between transport-induced amplification and dissipation. At early times, the growth of $A(\lambda,t)$ dominates, leading to an increase of the mismatch contribution. At later times, exponential dissipation takes over, forcing eventual decay.
Sufficient condition for double descent.
We therefore obtain the following sufficient condition: if there exists a time interval and a non-negligible spectral region
$\lambda\in[\lambda_{-},\lambda_{+}]\quad\text{such that}\quad\partial_{t}A(\lambda,t)>0,$
then the test loss $L_{\mathrm{te}}(t)$ is necessarily non-monotone, even though the training loss $L_{\mathrm{tr}}(t)$ decreases monotonically. This non-monotonicity manifests as the classical double-descent curve.
Importantly, this mechanism is entirely in-domain. It does not rely on distribution shift or adversarial effects, but follows generically from the interaction between renormalizable spectral transport and finite-sample spectral mismatch.